# Gaussian in the Wild: 3D Gaussian Splatting for Unconstrained Image Collections

### v1

@@ Abstract
Novel view synthesis from unconstrained in-the-wild images remains a meaningful but challenging task. The photometric variation and transient occluders in those unconstrained images make it difficult to reconstruct the original scene accurately. Previous approaches tackle the problem by introducing a global appearance feature in Neural Radiance Fields (NeRF). However, in the real world, the unique appearance of each tiny point in a scene is determined by its independent intrinsic material attributes and the varying environmental impacts it receives. Inspired by this fact, we propose Gaussian in the wild (GS-W), a method that uses 3D Gaussian points to reconstruct the scene and introduces separated intrinsic and dynamic appearance feature for each point, capturing the unchanged scene appearance along with dynamic variation like illumination and weather. Additionally, an adaptive sampling strategy is presented to allow each Gaussian point to focus on the local and detailed information more effectively. We also reduce the impact of transient occluders using a 2D visibility map. More experiments have demonstrated better reconstruction quality and details of GS-W compared to previous methods, with a 1000×1000$\times$1000 × increase in rendering speed. Video results and code are available at https://eastbeanzhang.github.io/GS-W/.
@@
Figure 1: With an unconstrained image collection input, GS-W can render novel views with appearance tuning, achieving state-of-the-art quality and 1000$\times$ rendering speed.
@@ 1 Introduction
  • Experimental results demonstrate that our method not only outperforms the previous state-of-the-art methods in terms of quality but also surpasses them in rendering speed by over 1000×1000$\times$1000 ×.
@@ 2 Related Work
Diverse 3D representations are developed to represent the geometric and appearance information of three-dimensional objects or scenes, among which implicit and explicit representation are two common methods for practical applications like 3D object generation and scene reconstruction. Implicit representation takes a coordinate as input and outputs corresponding scene information, like occupancy[20], distance[25], color, and density. NeRF [22] is an outstanding work among them which models the object information of each spatial point as radiation and volume density. NeRF indicates that using MLP can represent complex scenes and render novel photo-realistic views with the help of volume rendering. On the contrary, explicit representation describes the scene by optimizing a 3D model constructed by meshes[40, 13, 14], point clouds[27, 28, 35], voxels[42, 43, 33] and so on. More recently, methods represented by 3DGS[15] have entered researchers’ vision by their real-time rendering speed while preserving high-resolution synthesis quality. An efficient tile-based rasterizer is introduced by 3DGS to splat Gaussians to the image plane and accelerate the rendering speed many times compared to previous methods. Many researchers are exploring its potential by extending it to various tasks[41, 29, 47]. Several hybrid representations [4, 3, 5, 9, 34] are also emerging and creating more possibilities.
@@ 2.2 Novel view synthesis
Synthesizing arbitrary views of a scene using a set of 2D images is a long-standing problem in computer vision. Many NeRF-based methods[23, 10, 48, 11, 30] have achieved expressive synthesis quality [1, 44, 37] along with good view-consistency[7, 38, 24]. Mip-NeRF[1] replaces the rays with a 3D conical frustum and proposes integrated position embedding to anti-aliasing. Instant-NGP [23] introduces multi-resolution hash encoding that permits the use of a smaller network to reduce the training cost. There are also some Gaussian-based methods[46, 18, 8, 49] contributing to this task by modifying 3DGS. For example, to handle scenes with specular elements, Spec-Gaussian[46] departs from using spherical harmonics and instead adopts an anisotropic spherical Gaussian appearance field to model each point. Since these aforementioned methods all assume that input images are captured in a static scene, their performance declines intensely when reconstructing from unconstrained photo collections. Thus, several attempts[21, 19, 6, 45, 17] are proposed to address this challenging in-the-wild task by handling appearance variation and transient occluders. This field is still developing with some remaining issues and look forward to more contributing work.
As one of them, our proposed method tries to push the field one step forward by achieving a more delicate and flexible synthesis with an acceleration in rendering speed, through our modifications mentioned in Sec. 4.
@@ 4 Method
4.1.1 Appearance.
Since 3DGS[15] is designed for reconstructing static scenes, GS-W abandons the conventional color modeling approach using spherical harmonic coefficients. Instead, in the subsequent section, we introduce a new appearance feature $af_{i}$ for each Gaussian point, adapting to variations in the reference image by fusing intrinsic appearance feature $sf_{i}$ with dynamic appearance feature $df_{i}$ extracted from the image.
4.1.2 Transient object.
Handling transient objects is also challenging for 3DGS, in which Gaussian points around transient object regions may receive gradients and move or grow, bringing the emergence of meaningless floating points and rendering artifacts. Therefore, a visibility map is employed to mitigate this issue in Sec. 4.4. In addition, we maintain the pruning and growing strategies for points to reconstruct buildings that are absent in the initial point cloud.
@@ 4.2 Dynamic appearance features modeling
4.2.1 Projection feature map.
We extract 2D features from the image $I_{gt}$ and then map each point to the 2D space for feature sampling. With the known camera pose of the image, the mapping relationship from 3D points to the 2D image can be determined. Therefore, from the image we extract a projection feature map $F^{P}$, onto which each 3D point is projected using a projection matrix $P$, followed by the bilinear interpolation feature sampling, as illustrated below:
@@
4.2.2 K feature maps.
Limited by the single-view reference image $I_{gt}$, features sampled from the projection feature map are identical along the same ray, which is unreasonable under uneven lighting. Besides, since the reference image sometimes only contains part of the scene, many points may not be projected into effective regions for obtaining valid feature samples, causing inconsistencies in new viewpoints. To address these issues and grant Gaussian points the freedom to focus on diverse information, we propose extracting additional $K$ feature maps $(F^{1},F^{2}...F^{K})$ from $I_{gt}$ to construct a high-dimensional sampling space. Each Gaussian point is mapped to these maps respectively for adaptive sampling, allowing them to focus on high-frequency features.
4.2.3 Adaptive sampling.
It is necessary to map the Gaussian points to different positions on K feature maps, to better focus on diverse information. Inspired by the motive of each Gaussian point independently learning its own attributes, we consider it an efficient way to allow every point to determine its sampling positions through self-learning. Therefore, we assign each Gaussian point with K learnable sampling coordinate attributes $(sc^{1}_{i},sc^{2}_{i}...sc^{K}_{i})$, enabling them to adaptively select the information they need to focus on. The sampling process using these K sampling coordinates in K feature maps is as follows:
@@
4.2.4 Feature maps extraction.
We utilize a Unet[31] model with ResNet[12] backbone to generate both the projection feature map and K feature maps. The Unet model takes the reference image $I_{gt}\in\mathbb{R}^{3\times H\times W}$ as input and produces a 2D feature map $F\in\mathbb{R}^{16(K+1)\times H\times W}$ of the same spatial size. Subsequently, this feature map is evenly divided along the channel dimension into $(K+1)$ feature maps, each serving as one of the K feature maps $(F^{1},F^{2}...F^{K})$, and the projection feature map $F^{P}$. We choose Unet as the feature extractor due to its simplicity and effectiveness in extracting features from images at the same scale.
@@ 4.3 Intrinsic and dynamic appearance
4.3.1 Separation of intrinsic and dynamic appearance.
As mentioned before, an object’s appearance is influenced by both its intrinsic material as well as surface properties, and dynamic environmental factors. However, previous methods like Ha-NeRF primarily rely on image features, positional data, and viewing direction to decode appearance using MLPs. This implicit modeling of the scene’s intrinsic attributes through MLPs makes it difficult to express the high-frequency features solely with small MLPs, thus relying on the extracted dynamic features to capture comprehensive information. Consequently, such blurring of features hinders the model’s ability to distinguish between intrinsic and dynamic appearance accurately, particularly in scenarios involving changes in illumination and weather conditions. To address this, we explicitly separate the scene appearance into two forms: intrinsic and dynamic appearance features.
@@
4.3.2 Fusion of intrinsic and dynamic appearance features.
After acquiring independent dynamic appearance features for each Gaussian point through adaptive sampling, it’s essential to combine them with the corresponding intrinsic appearance features to generate a comprehensive appearance feature $af_{i}$. To accomplish this, we design a fusion network $M_{f}$ composed of two MLPs, which takes two appearance features and position information as inputs and produces a holistic appearance feature influenced by both. Specifically:
@@ 5.1 Implementation details
Table 1: Quantitative results on three datasets. The bold and the underline represent the best and second-best results, respectively. Our method outperforms the previous methods across all datasets on PSNR, SSIM, and LPIPS.
@@
Table 2: Comparison of rendering speed with one RTX 3090 GPU on three datasets with a resolution of 800 $\times$ 800, measured by FPS(Frames Per Second). Ours-cache means the appearance feature is cached for each point when synthesizing novel views.
@@ 5.2 Evaluation
5.2.1 Dataset, metrics, baseline.
We evaluate our proposed method on three datasets: Brandenburg Gate, Sacre Coeur, and Trevi Fountain, which include varying appearances and transient objects. For quantitative comparison, we use PSNR, SSIM[39], and LPIPS[50] as metrics to assess the performance of our method. We also present rendered images generated from the same pose as the input view for visual inspection. To demonstrate the superiority of our method, we evaluate our proposed method against 3DGS[15], NeRF-W [19], Ha-NeRF[6], and CR-NeRF[45].
5.2.2 Quantitative comparison.
Quantitative results are shown in Tab. 1. 3DGS performs poorly on both PSNR and SSIM metrics, as it does not explicitly model the appearance variation and transient objects. NeRF-W and Ha-NeRF achieve moderate performance by introducing a global appearance embedding and anti-transient module. It’s worth noting that NeRF-W needs to optimize the appearance embedding of test images, thus the comparison with NeRF-W is unfair. CR-NeRF achieves competitive performance due to cross-ray manner. Utilizing an adaptive sampling strategy that allows each point to focus on local details, our method outperforms the baselines on three datasets in terms of PSNR, SSIM, and LPIPS, which verifies that we can capture more details and render higher-quality images.
5.2.3 Render speed.
To compare the render speed of different methods during inference, we experiment on the three datasets by setting the image resolution to $800\times 800$ and calculating the average rendering time per image using a single RTX 3090 GPU. The time taken for feature extraction from the reference images in Ha-NeRF, CR-NeRF, and our method is included in the overall inference time. As shown in Tab. 2, our method achieves a significant improvement in rendering speed, which is 1000$\times$ faster than previous NeRF-based methods.
Figure 3: Qualitative experimental results on three unconstrained datasets. GS-W recovers finer details of appearance(e.g. the horse sculpture in Brandenburg, the sky and clouds in Sacre, the light on columns, and the color of windows in Trevi). Moreover, GS-W reconstructs more consistent and detailed scenes (e.g. the distant tower in Brandenburg, the cavities in Sacre, and the distant building in Trevi).
@@
5.2.4 Qualitative comparison.
Fig. 3 presents the qualitative result for all methods. NeRF-W and Ha-NeRF can model varying appearances from the reference image by introducing a global appearance embedding. CR-NeRF can reconstruct better geometry and model appearance variation compared with Ha-NeRF and NeRF-W. However, they all struggle to reconstruct the details of scenes in the distance and the intricate textures of the scenes, e.g. the door pillars and distant tower in Brandenburg, the cavity in Sacre, and the distant buildings in Trevi. In contrast, thanks to the higher-frequency dynamic appearance features, our method recovers more accurate appearance details, e.g. the horse sculpture in Brandenburg, the sky and clouds in Sacre, and the light on columns in Trevi.
@@ 5.3 Ablation studies
We summarize the ablation studies of our method on Brandenburg, Sacre, and Trevi datasets in Tab. 3 and produce qualitative results in Fig. 4 to validate the effectiveness of each component.
5.3.1 Without the visibility map.
Removing the transient object handling module leads to higher metric performances but introduces artifacts in the rendered image due to the influence of transient objects, as depicted in Fig. 4. Since most test images lack dynamic objects, these artifacts may not significantly impact the metrics.
5.3.2 Without K feature maps or Projection feature map.
The removal of K feature maps or the projection feature map results in performance degradation. Especially, synthesizing novel views without the K feature maps not only reduces the ability to capture information from the reference image but also produces view-inconsistent appearances, as shown in Fig. 4.
5.3.3 Without adaptive sampling.
Retaining K feature maps while keeping the sampling coordinates fixed for each point significantly decreases performance. This highlights the importance of our adaptive sampling strategy for K feature maps, enabling Gaussian points to adaptively focus on local and detailed features.
Table 3: Ablation studies on three datasets. The bold and the underline represent the best and second-best results, respectively. See Sec. 5.3 for detailed descriptions.
@@
full (Ours) 27.96 0.9319 0.0862 23.24 0.8632 0.1300 22.91 0.0814 0.1563
Figure 4: Ablation studies by visualization. The first row represents the reference images and the corresponding rows represent the rendered images from a novel view. With K feature maps and a visibility map, our method is capable of performing view-consistent appearance and reducing artifacts.
5.3.4 Without separation.
We eliminate the intrinsic feature and predict the color of Gaussian points by the dynamic appearance feature only. The results show a noticeable decrease in both LPIPS and SSIM, which are more in line with human visual perception. This illustrates the differentiation between the two is beneficial for the model to comprehend and recover the appearance accurately.

### v2

@@ Abstract
Novel view synthesis from unconstrained in-the-wild images remains a meaningful but challenging task. The photometric variation and transient occluders in those unconstrained images make it difficult to reconstruct the original scene accurately. Previous approaches tackle the problem by introducing a global appearance feature in Neural Radiance Fields (NeRF). However, in the real world, the unique appearance of each tiny point in a scene is determined by its independent intrinsic material attributes and the varying environmental impacts it receives. Inspired by this fact, we propose Gaussian in the wild (GS-W), a method that uses 3D Gaussian points to reconstruct the scene and introduces separated intrinsic and dynamic appearance feature for each point, capturing the unchanged scene appearance along with dynamic variation like illumination and weather. Additionally, an adaptive sampling strategy is presented to allow each Gaussian point to focus on the local and detailed information more effectively. We also reduce the impact of transient occluders using a 2D visibility map. More experiments have demonstrated better reconstruction quality and details of GS-W compared to NeRF-based methods, with a faster rendering speed. Video results and code are available at https://eastbeanzhang.github.io/GS-W/.
@@
Figure 1: With an unconstrained image collection input, GS-W can render novel views with appearance tuning, achieving state-of-the-art quality and faster rendering speed.
@@ 1 Introduction
  • Experimental results demonstrate that our method not only outperforms the state-of-the-art NeRF-based methods in terms of quality but also surpasses them in rendering speed by over 1000×1000$\times$1000 ×.
@@ 2 Related Work
Diverse 3D representations are developed to represent the geometric and appearance information of three-dimensional objects or scenes, among which implicit and explicit representation are two common methods for practical applications like 3D object generation and scene reconstruction. Implicit representation represents 3D data as continuous functions or fields, like occupancy fields[24], distance fields [29], color, and density. NeRF [26] is an outstanding work among them which models the scene as a continuous field of density and radiance. NeRF indicates that using MLP can represent complex scenes and render novel photo-realistic views with the help of volume rendering. On the contrary, explicit representation describes scenes by storing and manipulating 3D data using discrete structures like meshes[45, 14, 15], point clouds[31, 32, 40], voxels[47, 48, 38] and so on. More recently, methods represented by 3DGS[16] have entered researchers’ vision by their real-time rendering speed while preserving high-resolution synthesis quality. An efficient tile-based rasterizer is introduced by 3DGS to splat Gaussians to the image plane and accelerate the rendering speed many times compared to previous methods. Many researchers are exploring its potential by extending it to various tasks[46, 33, 52]. Several hybrid representations [4, 3, 5, 10, 39] are also emerging and creating more possibilities.
@@ 2.2 Novel view synthesis
Synthesizing arbitrary views of a scene using a set of 2D images is a long-standing problem in computer vision. Many NeRF-based methods[27, 11, 53, 12, 34] have achieved expressive synthesis quality [1, 49, 42] along with good view-consistency[7, 43, 28]. Mip-NeRF[1] replaces the rays with a 3D conical frustum and proposes integrated position embedding to anti-aliasing. Instant-NGP [27] introduces multi-resolution hash encoding that permits the use of a smaller network to reduce the training cost. There are also some Gaussian-based methods[51, 22, 9, 54] contributing to this task by modifying 3DGS. For example, to handle scenes with specular elements, Spec-Gaussian[51] departs from using spherical harmonics and instead adopts an anisotropic spherical Gaussian appearance field to model each point. Since these aforementioned methods all assume that input images are captured in a static scene, their performance declines intensely when reconstructing from unconstrained photo collections. Thus, several attempts[25, 23, 36, 6, 50, 19] are proposed to address this challenging in-the-wild task by handling appearance variation and transient occluders. Other works [20, 21] focus on scenes with time-varying appearances, while methods [55, 18, 8] use physical rendering models for diverse lighting conditions. This field still faces some remaining issues and looks forward to advancements.
As one of them, our proposed method tries to push the field one step forward by achieving a more delicate and flexible synthesis with higher rendering speed, through our modifications mentioned in Sec. 4.
@@ 4 Method
Appearance. Since 3DGS[16] is designed for reconstructing static scenes, GS-W abandons the conventional color modeling approach using spherical harmonic coefficients. Instead, in the subsequent section, we introduce a new appearance feature $af_{i}$ for each Gaussian point, adapting to variations in the reference image by fusing intrinsic appearance feature $sf_{i}$ with dynamic appearance feature $df_{i}$ extracted from the image.
Transient object. Handling transient objects is also challenging for 3DGS, in which Gaussian points around transient object regions may receive gradients and move or grow, bringing the emergence of meaningless floating points and rendering artifacts. Therefore, a visibility map is employed to mitigate this issue in Sec. 4.4. In addition, we maintain the pruning and growing strategies for points to reconstruct buildings that are absent in the initial point cloud.
@@ 4.2 Dynamic appearance features modeling
Projection feature map. We extract 2D features from the image $I_{gt}$ and then map each point to the 2D space for feature sampling. With the known camera pose of the image, the mapping relationship from 3D points to the 2D image can be determined. Therefore, from the image we extract a projection feature map $F^{P}$, onto which each 3D point is projected using a projection matrix $P$, followed by the bilinear interpolation feature sampling, as illustrated below:
@@
K feature maps. Limited by the single-view reference image $I_{gt}$, features sampled from the projection feature map are identical along the same ray, which is unreasonable under uneven lighting. Besides, since the reference image sometimes only contains part of the scene, many points may not be projected into effective regions for obtaining valid feature samples, causing inconsistencies in new viewpoints. To address these issues and grant Gaussian points the freedom to focus on diverse information, we propose extracting additional $K$ feature maps $(F^{1},F^{2}...F^{K})$ from $I_{gt}$ to construct a high-dimensional sampling space. Each Gaussian point is mapped to these maps respectively for adaptive sampling, allowing them to focus on high-frequency features.
Adaptive sampling. It is necessary to map the Gaussian points to different positions on K feature maps, to better focus on diverse information. Inspired by the motive of each Gaussian point independently learning its own attributes, we consider it an efficient way to allow every point to determine its sampling positions through self-learning. Therefore, we assign each Gaussian point with K learnable sampling coordinate attributes $(sc^{1}_{i},sc^{2}_{i}...sc^{K}_{i})$, enabling them to adaptively select the information they need to focus on. The sampling process using these K sampling coordinates in K feature maps is as follows:
@@
Feature maps extraction. We utilize a Unet[35] model with ResNet[13] backbone to generate both the projection feature map and K feature maps. The Unet model takes the reference image $I_{gt}\in\mathbb{R}^{3\times H\times W}$ as input and produces a 2D feature map $F\in\mathbb{R}^{16(K+1)\times H\times W}$ of the same spatial size. Subsequently, this feature map is evenly divided along the channel dimension into $(K+1)$ feature maps, each serving as one of the K feature maps $(F^{1},F^{2}...F^{K})$, and the projection feature map $F^{P}$. We choose Unet as the feature extractor due to its simplicity and effectiveness in extracting features from images at the same scale.
@@ 4.3 Intrinsic and dynamic appearance
Separation of intrinsic and dynamic appearance. As mentioned before, an object’s appearance is influenced by both its intrinsic material as well as surface properties, and dynamic environmental factors. However, previous methods like Ha-NeRF primarily rely on image features, positional data, and viewing direction to decode appearance using MLPs. This implicit modeling of the scene’s intrinsic attributes through MLPs makes it difficult to express the high-frequency features solely with small MLPs, thus relying on the extracted dynamic features to capture comprehensive information. Consequently, such blurring of features hinders the model’s ability to distinguish between intrinsic and dynamic appearance accurately, particularly in scenarios involving changes in illumination and weather conditions. To address this, we explicitly separate the scene appearance into two forms: intrinsic and dynamic appearance features.
@@
Fusion of intrinsic and dynamic appearance features. After acquiring independent dynamic appearance features for each Gaussian point through adaptive sampling, it’s essential to combine them with the corresponding intrinsic appearance features to generate a comprehensive appearance feature $af_{i}$. To accomplish this, we design a fusion network $M_{f}$ composed of two MLPs, which takes two appearance features and position information as inputs and produces a holistic appearance feature influenced by both. Specifically:
@@ 5.1 Implementation details
Table 1: Quantitative results on the test set of three PhotoTourism scenes. The bold and the underline represent the best and second-best results, respectively. Our method outperforms the previous methods across all scenes on PSNR, SSIM, and LPIPS.
@@
Table 2: Comparison of rendering speed with one RTX 3090 GPU on three scenes with a resolution of 800 $\times$ 800, measured by FPS(Frames Per Second). Ours-cache means the appearance feature is cached for each point when synthesizing novel views.
@@ 5.2 Evaluation
Dataset, metrics, baseline. We evaluate our proposed method on three scenes from the PhotoTourism dataset: Brandenburg Gate, Sacre Coeur, and Trevi Fountain, which include varying appearances and transient objects. For quantitative comparison, we use PSNR, SSIM[44], and LPIPS[56] as metrics to assess the performance of our method. We also present rendered images generated from the same pose as the input view for visual inspection. To demonstrate the superiority of our method, we evaluate our proposed method against 3DGS[16], NeRF-W [23], Ha-NeRF[6], and CR-NeRF[50].
Quantitative comparison. Quantitative results are shown in Tab. 1. 3DGS performs poorly on both PSNR and SSIM metrics, as it does not explicitly model the appearance variation and transient objects. NeRF-W and Ha-NeRF achieve moderate performance by introducing a global appearance embedding and anti-transient module. It’s worth noting that NeRF-W needs to optimize the appearance embedding of test images, thus the comparison with NeRF-W is unfair. CR-NeRF achieves competitive performance due to cross-ray manner. Utilizing an adaptive sampling strategy that allows each point to focus on local details, our method outperforms the baselines on three scenes in terms of PSNR, SSIM, and LPIPS, which verifies that we can capture more details and render higher-quality images.
Render speed. To compare the render speed of different methods during inference, we experiment on the three scenes by setting the image resolution to $800\times 800$ and calculating the average rendering time per image using a single RTX 3090 GPU. The time taken for feature extraction from the reference images in Ha-NeRF, CR-NeRF, and our method is included in the overall inference time. As shown in Tab. 2, our method achieves a significant improvement in rendering speed, which is 1000$\times$ faster than previous NeRF-based methods.
Figure 3: Qualitative results on the test set of three PhotoTourism scenes. GS-W recovers finer details of appearance(e.g. the horse sculpture in Brandenburg, the sky and clouds in Sacre, the light on columns, and the color of windows in Trevi). Moreover, GS-W reconstructs more consistent and detailed scenes (e.g. the distant tower in Brandenburg, the cavities in Sacre, and the distant building in Trevi).
@@
Qualitative comparison. Fig. 3 presents the qualitative result for all methods. NeRF-W and Ha-NeRF can model varying appearances from the reference image by introducing a global appearance embedding. CR-NeRF can reconstruct better geometry and model appearance variation compared with Ha-NeRF and NeRF-W. However, they all struggle to reconstruct the details of scenes in the distance and the intricate textures of the scenes, e.g. the door pillars and distant tower in Brandenburg, the cavity in Sacre, and the distant buildings in Trevi. In contrast, thanks to the higher-frequency dynamic appearance features, our method recovers more accurate appearance details, e.g. the horse sculpture in Brandenburg, the sky and clouds in Sacre, and the light on columns in Trevi.
@@ 5.3 Ablation studies
We summarize the ablation studies of our method on Brandenburg, Sacre, and Trevi datasets in Tab. 3 and produce qualitative results in Fig. 4 and Fig. 5 to validate the effectiveness of each component.
Without the visibility map. Removing the transient object handling module leads to higher metric performances but introduces artifacts in the rendered image due to the influence of transient objects, as depicted in Fig. 4. Since most test images lack dynamic objects, these artifacts may not significantly impact the metrics.
Without K feature maps or projection feature map. The removal of K feature maps or the projection feature map results in performance degradation. Especially, synthesizing novel views without the K feature maps not only reduces the ability to capture information from the reference image but also produces view-inconsistent appearances, as shown in Fig. 4.
Without adaptive sampling. Retaining K feature maps while keeping the sampling coordinates fixed for each point significantly decreases performance. This highlights the importance of our adaptive sampling strategy for K feature maps, enabling Gaussian points to adaptively focus on local and detailed features.
Table 3: Ablation studies on three scenes. The bold and the underline represent the best and second-best results, respectively. See Sec. 5.3 for detailed descriptions.
@@
full (Ours) 27.96 0.9319 0.0862 23.24 0.8632 0.1300 22.91 0.8014 0.1563
Figure 4: Ablation studies by visualization. The first row represents the reference images and the corresponding rows represent the rendered images from a novel view. Our full method is capable of performing view-consistent appearance and reducing artifacts.
Without separation. We eliminate the intrinsic feature and predict the color of Gaussian points by the dynamic appearance feature only. The results in Tab. 3 show a noticeable decrease in both LPIPS and SSIM, which are more in line with human visual perception. In Fig. 5, the absence of dynamic appearance features results in incomplete scene appearance. This emphasizes the significance of intrinsic feature in retaining essential scene characteristics. Both illustrate that separation is beneficial for the model to accurately comprehend and reconstruct sharp appearances.